OpenStudy (anonymous):

Which ratio forms a proportion with 12/10

OpenStudy (unklerhaukus):

\[\frac {12}{10}\]factorize each number \[\frac {2\times6}{2\times5}\]now cancel common factors,

OpenStudy (anonymous):

uhh..

OpenStudy (anonymous):

thank u

OpenStudy (unklerhaukus):

\[\frac {2\times6}{2\times5}\]\[=\frac {\cancel{2\times}6}{\cancel{2\times}5}\]\[=\frac 65\]